OpenStudy (anonymous):

factorise x^2 - y^2 - 2x + 6y - 8 Pls show steps foe getting the solution...

OpenStudy (nikvist):

\[x^2-y^2-2x+6y-8=x^2-2x+1-y^2+6y-9=\]\[=(x-1)^2-(y-3)^2=(x-y+2)(x+y-4)\]

OpenStudy (anonymous):

Thanks a ton nikvist !!!!! U deserve more than one medal but i can giv only one....☺☻